# StarStream on Gaia: Stream discovery and mass loss rate of globular clusters

### v1

@@ 1 Introduction
By applying variations of the matched-filter technique (e.g., C. M. Rockosi et al., 2002; C. J. Grillmair, 2009; E. J. Bernard et al., 2014; N. Shipp et al., 2018) to both the color–magnitude diagram (CMD) and proper motion space, astronomers have identified over one hundred thin, elongated structures recognized as stellar streams (see the review by A. Bonaca & A. M. Price-Whelan, 2025). Notably, R. Ibata et al. (2024) employed the STREAMFINDER algorithm (K. Malhan & R. A. Ibata, 2018) on Gaia DR3 (Gaia Collaboration et al., 2023) to uncover 87 thin streams, while A. Hallin et al. (2025) combined the Via Machinae method (D. Shih et al., 2021) with the Cathode algorithm (A. Hallin et al., 2022) to detect around 80 thin streams in Gaia DR2.
@@
The high detection quality makes StarStream a powerful tool to uncover GC streams that may have been missed by previous methods. Its quantified performance further allows us to derive unbiased estimates of the mass loss rate for these clusters. In this work, we apply StarStream to Gaia DR3 fields around MW GCs. In §2, we provide an overview of StarStream and the Gaia DR3 dataset. We then present the discovery of new streams in §3, followed by the calculation of the mass loss rates in §4. Finally, we summarize and discuss our findings in §5.
2 Method
@@ 2.1 Overview of StarStream
We release tracer particles over the last 1 Gyr assuming a uniform ejection rate $\dot{N}_{\rm tracer}=4\ {\rm Myr}^{-1}$. We then sample the mass of each particle from the P. Kroupa (2001) initial mass function, with the minimum stellar mass being set to the lowest possible mass of the closest tracer particle that remains above the detection limit. Based on the stellar mass, we calculate the color and magnitude of each tracer particle from the MESA Isochrones and Stellar Tracks (MIST, A. Dotter, 2016; U. Meštrić et al., 2022) model. In C25, we have verified that these settings are sufficient to fully sample the position space, the proper motion space, and the CMD.
The KDE is constructed on these tracers with no correlation between each dimension. We set the kernel bandwidth to 0.1 times the standard deviation of all tracer particles in each dimension, except that we fix the bandwidth to 0.1 for magnitudes and 0.02 for colors. Varying these values by a factor of 0.520.5-2 has a negligible effect on our results. It is worth noting that we convolve the kernels with the observational uncertainties when evaluating the KDE for observed stars. This is equivalent to increasing the bandwidths of KDE kernels.
@@
2.2 Observational datasets
@@
To obtain the isochrone of each GC in the CMD, we use the MIST model, which requires the age, metallicity, and color excess E(BV)E(B-V) due to extinction assuming the J. A. Cardelli et al. (1989) reddening law. We use E(BV)E(B-V) from the measurements by D. Massari et al. [^3] Private comm. in the context of the Cluster Ages to Reconstruct the Milky Way Assembly (CARMA) project (D. Massari et al., 2023)., or from the 2010 edition of the catalog by W. E. Harris (1996) [^4] https://physics.mcmaster.ca/~harris/mwgc.dat for GCs not included in the former. For the remaining streams without reliable extinction measurements, we use the extinction map by D. J. Schlegel et al. (1998) recalibrated by E. F. Schlafly & D. P. Finkbeiner (2011). The three sources of E(BV)E(B-V) typically differ only by $\lesssim 30\%$ for the overleaping GCs. To avoid extreme extinction that affects our detection quality, we exclude $\sim 20\%$ GCs with E(BV)>1E(B-V)>1, leaving 128 GCs in our final sample.
@@
3 Detecting streams in the Milky Way
3.1 Method validation
@@
We note that these metrics strongly depend on the progenitor’s extinction $A_{V}$ and background density. The latter is characterized by the number of background stars $N_{\rm bg}$ (accounting for the CMD selection in §2.2) within the $10^{\circ}$ search radius. In the upper row of Fig. 1, we show the completeness and purity as functions of these metrics. The values are directly taken from the with-extinction case of C25 (see §3.7 therein), excluding extremely high extinction cases following §2.2. As expected, both completeness and purity decrease with $A_{V}$ and $N_{\rm bg}$. They become less than 20% at $A_{V}>0.6$ or $N_{\rm bg}>6\times 10^{6}$. These values approximately correspond to the low-Galactic latitude region $|b|\lesssim 20^{\circ}$, where the high dust reddening and background contamination significantly affect the detection quality. On the other hand, the method achieves acceptable completeness and purity for the low-$A_{V}$ and low-$N_{\rm bg}$ GCs. For mock GCs with $A_{V}<0.6$ and $N_{\rm bg}<6\times 10^{6}$, the median completeness and purity are 50% and 59%, respectively. These values can even grow to $60\%-80\%$ near the Galactic poles.
@@ 3.2 New streams
In Table 1, we list the key properties of the 34 streams in our high-quality sample. We provide information about the low-quality sample in Appendix B, Table 2. We also plot the high-quality sample in the great circle frame $\phi_{1}$$\phi_{2}$ in Fig. 2. The $\phi_{1}$$\phi_{2}$ frame is defined such that the progenitor GC is at $(0,0)$, and the stream is elongated along $\phi_{1}$. We achieve this by rotating the sky coordinates to minimize the standard deviation of $\phi_{2}$ for simulated tracer particles. We rank the streams in Fig. 2 by the length of simulated streams, which is characterized by the 90th percentile of angular separation, $r_{90}$. Since we only detect streams within the $10^{\circ}$ search radius around the GC, there are 17 streams with $r_{90}$ exceeding this radius. In §3.3, we perform a followup detection for these streams by extending the search radius to $20^{\circ}$.
@@
In Fig. 3, we show the CMD of detected stream stars using Gaia photometry. Most stars are gathered around or below the main sequence turn off. Since StarStream takes into account the color uncertainty that typically grows with the $G$-magnitude, the color spread of detected stars also becomes larger near the fainter end. On the other hand, the color spread of simulated streams is only due to the distance spread and is thus almost invariant of $G$-magnitude.
@@
Table 1: Summary of GC and stream properties for the high-quality sample. The Galactic longitude $l$, latitude $b$, heliocentric distance $d_{\odot}$, cluster mass $M_{\rm GC}$, and 3D half-mass radius $r_{\rm h}$ are taken from the M. Hilker et al. (2019) catalog. The extinction $A_{V}$ is either from D. Massari et al., W. E. Harris (1996) catalog, or D. J. Schlegel et al. (1998) map, see §$2.$ $N_{\rm bg}$ is the number of CMD-selected Gaia DR3 stars within the $10^{\circ}$ search radius. The number of detections $N_{\rm detect}$, tidal frequency $\Omega_{\rm tid}$, and cluster mass loss rate $|\dot{M}|$ are calculated in this work. A plain ASCII version is available at https://github.com/ybillchen/StarStream_DR.
@@
3.3 Extended detection for long streams
@@
In Fig. 4, we show the three streams, NGC 5272, 1851, and 5024, with more than 10 new detections outside $10^{\circ}$. Note that the number of background stars approximately triples in the annulus, significantly reducing the signal-to-noise ratio in this region. Therefore, we should not expect the completeness and purity to stay unchanged in the extended area.
@@
3.4 Comparison with previous detections
@@
Recently, P. B. Kuzma et al. (2025) identified potential tidal structures within $5^{\circ}$ around 22 MW GCs using the Pristine-Gaia synthetic catalog from Pristine Data Release 1 (E. Starkenburg et al., 2017; N. F. Martin et al., 2024). 13 of them overlap our high-quality sample: NGC 362, 1261, 1851, 1904, 5272, 6205, 6341, 6934, 6981, 7078, 7089, 7099. It is worth noting that many of their tidal structures do not show stream-like feature. This agrees with our finding that although many GC streams are “irregular”, the tidal streams around GCs are more common than was previously detected.
4 Mass loss rate of globular clusters
4.1 Calculation of mass loss rate
@@
Y. Chen et al. (2025b) approximates $f_{\rm sel}$ as a Gaussian tube around the stream track when analyzing the R. Ibata et al. (2024) stream catalog. Here, however, we can make further simplification considering the quantified detection ratio $f_{\rm detect}$ between the expectation of the number of detections $N_{\rm detect}$ by StarStream and the true number $N_{\rm true}$ (see C25, ). In Appendix A, we prove that simply replacing the integral in Eq. (1) by the total number of tracer particles $N_{\rm tracer}$ times $f_{\rm detect}$ yields an unbiased estimate of the mass loss rate:
@@
where $f_{\rm detect}\equiv N_{\rm detect}/N_{\rm true}$ varies with individual streams. Using the tests by C25, we find that $f_{\rm detect}$ is always centered at 0.9 for the high-quality sample, with a log-normal scatter that depends on the background density $N_{\rm bg}$. In Appendix A, we approximate the standard deviation of the log-normal scatter $\sigma_{\log f}$ as a linear function of $\log_{10}N_{\rm bg}$: $\sigma_{\log f}$ increases from 0.15 dex to 0.5 dex when $\log_{10}N_{\rm bg}$ increases from 5.5 to 7, and is fixed to $\sigma_{\log f}=0.15$ dex below $\log_{10}N_{\rm bg}<5.5$ (see Fig. 6). Since the purity and completeness of the low-quality sample are both low with large variation, the scatter can be more than 1 dex. Therefore, we exclude the low-quality sample for the calculation of the mass loss.
Note that $M_{\rm sel}$ and $N_{\rm tracer}$ only account for tracer particles within the $10^{\circ}$ search radius. The detection ratio is not calibrated outside $10^{\circ}$ and is likely much lower. We thus exclude extended stream segments outside this radius by §3.3 for the subsequent calculation.
@@
Since we focus on the $10^{\circ}$ cone around each GC, the heliocentric distance of each stream member $d_{\odot,i}$ is very close to the distance to the center GC, $d_{\odot,\rm GC}$. Using the simulated stream, we verify that the standard deviation of distances is typically $0.025$ dex (at most $0.08$ dex). Therefore, the distance spread contributes negligibly to the total uncertainty since the intrinsic scatter due to the detection method itself is already >0.15>0.15 dex (see Appendix A). It is thus reasonable to adopt $d_{\odot,i}\approx d_{\odot,\rm GC}$. Correspondingly, we can define $m_{{\rm min},i}=m_{\rm min}(d_{\odot,\rm GC})\equiv m_{\rm min,GC}$ and $w_{i}=w(m_{\rm min,GC})\equiv w_{\rm GC}$. Eq. (3) is thus simplified to
@@
4.2 Mass loss rate vs. other cluster properties
@@
Figure 5 shows the measured mass loss rate for the high-quality sample as functions of GC’s mass $M_{\rm GC}$, effective tidal frequency $\Omega_{\rm tid}$, and 3D half-mass radius $r_{\rm h}$. The values of $M_{\rm GC}$ and $r_{\rm h}$ are directly taken from the M. Hilker et al. (2019) catalog, while $\Omega_{\rm tid}$ is approximated by $\sqrt{2}$ times the orbital frequency. This definition is consistent with M. Gieles & O. Y. Gnedin (2023), who used the singular isothermal sphere profile and adopted F. Renaud et al. (2011)’s definition based on eigenvalues of the tidal tensor. The orbital frequency is computed via integrating the GC’s orbit in our Galactic potential model. These values are listed in Table 1.
@@
$\begin{array}[]{rl}|\dot{M}_{\rm ref}|&=7.4^{+1.5}_{-1.3}\,{\rm M}_{\odot}\,{\rm Myr^{-1}}\\ \sigma_{\rm int}&=0.28\pm 0.06\ {\rm dex}\\ a&=0.15\pm 0.12\\ c&=0.58\pm 0.34\end{array}$ (6)
@@
where the uncertainties are obtained from 1000 bootstrap resampling. We find that the $\dot{M}$$M_{\rm GC}$ slope $a$ is above but still consistent with zero within 1.3 standard deviations, and the value lies between the no black holes (BHs) model ($a=1/3$) and BHs models (a=1/3a=-1/3, assuming the initial mass $M_{\rm i}=2\times 10^{5}\,{\rm M}_{\odot}$) of M. Gieles & O. Y. Gnedin (2023). Our $a$ is lower than that in Y. Chen et al. (2025b, $a=0.66\pm 0.37$) by $1.4\sigma$. The $\dot{M}$$r_{\rm h}$ slope is above zero by $1.7\sigma$, still being consistent with Y. Chen et al. (2025b, $c=0.12\pm 0.72$). We also observe larger intrinsic scatter compared to Y. Chen et al. (2025b). This is likely because our measurement includes the uncertainty of the detection method itself, whereas Y. Chen et al. (2025b) assumed zero uncertainty for the detection method. Although trends can differ, the amplitude of $\dot{M}$ in this work is consistent with M. Gieles & O. Y. Gnedin (2023) and Y. Chen et al. (2025b) within the intrinsic scatter across the full ranges of $M_{\rm GC}$, $\Omega_{\rm tid}$, and $r_{\rm h}$.
@@
In particular, we measured $|\dot{M}|\approx 10\ {\rm M_{\odot}\,Myr^{-1}}$ for Pal 5. This is higher than that in Y. Chen et al. (2025b, $|\dot{M}|\approx 3\ {\rm M_{\odot}\,Myr^{-1}}$) partially because of the 70% increase of $N_{\rm detect}$ (as mentioned in §3.4). However, this increase is insufficient to explain the 230% increase of $\dot{M}$. Since Y. Chen et al. (2025b) measured the average mass loss rate over the past $\sim 6$ Gyr for Pal 5 while this work only measures the past 1 Gyr, the different values likely suggest an accelerating mass loss rate for Pal 5. This scenario is consistent with the prediction by M. Gieles et al. (2021), who showed that Pal 5’s potential high BH abundance ($f_{\rm BH}\approx 20\%$) could accelerate the mass loss rate from the initial $5-10\ {\rm M_{\odot}\,Myr^{-1}}$ to $10-20\ {\rm M_{\odot}\,Myr^{-1}}$ near the end of its lifetime.
@@
The positive $\dot{M}$$r_{\rm h}$ correlation is likely dominated by the low-mass but high-$\dot{M}$ GCs similar to Pal 5, including NGC 7492, Pal 12, and Whiting 1. These GCs have $M_{\rm GC}\lesssim 2\times 10^{4}\,{\rm M}_{\odot}$ but half-mass radii above average, $r_{\rm h}\gtrsim 10$ pc, making them much “fluffier” than average GCs. Their high mass loss rates place them in closer agreement with M. Gieles & O. Y. Gnedin (2023)’s BHs models (a=1/3a=-1/3), while they are seemingly outlier of other models. M. Gieles et al. (2021) suggested that these GCs are also BH-rich and are reaching complete tidal dissolution. They also predicted higher-than-average mass loss rates for these GCs, $|\dot{M}|\approx 10\ {\rm M_{\odot}\,Myr^{-1}}$. Our measurements agree with their prediction, providing observational support for the BH-rich scenario.
5 Summary and discussion

### v2

@@ I Introduction
By applying variations of the matched-filter technique (e.g., C. M. Rockosi et al., 2002; C. J. Grillmair, 2009; E. J. Bernard et al., 2014; N. Shipp et al., 2018) to both the color–magnitude diagram (CMD) and proper motion space, astronomers have identified over one hundred thin, elongated structures recognized as stellar streams (see the review by A. Bonaca & A. M. Price-Whelan, 2025). Notably, R. Ibata et al. (2024) employed the STREAMFINDER algorithm (K. Malhan & R. A. Ibata, 2018) on Gaia DR3 (Gaia Collaboration et al., 2023) to uncover 87 thin streams, while A. Hallin et al. (2025) combined the Via Machinae method (D. Shih et al., 2021, 2024) with the Cathode algorithm (A. Hallin et al., 2022) to detect around 80 thin streams in Gaia DR2.
@@
The high detection quality makes StarStream a powerful tool to uncover GC streams that may have been missed by previous methods. Its quantified performance further allows us to derive unbiased estimates of the mass loss rate for these clusters. In this work, we apply StarStream to Gaia DR3 fields around MW GCs. In §II, we provide an overview of StarStream and the Gaia DR3 dataset. We then present the discovery of new streams in §III, followed by the calculation of the mass loss rates in §IV. Finally, we summarize and discuss our findings in §V.
II Method
@@ II.1 Overview of StarStream
We release tracer particles over the last 1 Gyr assuming a uniform ejection rate $\dot{N}_{\rm tracer}=4\ {\rm Myr}^{-1}$. We then sample the mass of each particle from the P. Kroupa (2001) initial mass function, with the minimum stellar mass being set to the lowest possible mass of the closest tracer particle that remains above the detection limit. Based on the stellar mass, we calculate the color and magnitude of each tracer particle from the MESA Isochrones and Stellar Tracks (MIST, A. Dotter, 2016; U. Meštrić et al., 2022) model. $\dot{N}_{\rm tracer}$ is a numerical parameter that controls only the number of tracers to sample position space, proper motion space, and the CMD. In C25, we verified that our adopted value is sufficient to fully sample these spaces.
The KDE is constructed on these tracers. The kernel is a product of one-dimensional kernels with the bandwidth adapted to 0.1 times the standard deviation of all tracer particles in each dimension, except that we fix the bandwidth to 0.1 for magnitudes and 0.02 for colors. Varying these values by a factor of 0.520.5-2 has a negligible effect on our results. It is worth noting that we convolve the kernels with the observational uncertainties when evaluating the KDE for observed stars. This is equivalent to increasing the bandwidths of KDE kernels. Our KDE is constructed in the full multidimensional space of spatial coordinates, proper motions, colors, and magnitudes, as illustrated in Fig. 1 of C25. Therefore, our stream model naturally captures the curved structure of streams across all of these dimensions.
@@
Although this method works for most streams, it assumes that there is no significant overlap between multiple streams. An exception was noted by Z. Yuan et al. (2020) for the GC pair NGC 5024 and 5053, which reside near the Galactic pole and are separated by only $\sim 1^{\circ}$. We therefore apply a special treatment using a three-component mixture model, as described in Appendix A.
II.2 Observational datasets
@@
To obtain the isochrone of each GC in the CMD, we use the MIST model, which requires the age, metallicity, and color excess E(BV)E(B-V) due to extinction assuming the J. A. Cardelli et al. (1989) reddening law. We use E(BV)E(B-V) from the measurements by D. Massari et al. [^3] Private comm. in the context of the Cluster Ages to Reconstruct the Milky Way Assembly (CARMA) project (D. Massari et al., 2023)., or from the 2010 edition of the catalog by W. E. Harris (1996) [^4] https://physics.mcmaster.ca/~harris/mwgc.dat for GCs not included in the former. For the remaining streams without reliable extinction measurements, we use the extinction map by D. J. Schlegel et al. (1998) recalibrated by E. F. Schlafly & D. P. Finkbeiner (2011). The three sources of E(BV)E(B-V) typically differ only by $\lesssim 30\%$ for the overlapping GCs. To avoid extreme extinction that affects our detection quality, we exclude $\sim 20\%$ GCs with E(BV)>1E(B-V)>1, leaving 128 GCs in our final sample.
@@
III Detecting streams in the Milky Way
III.1 Method validation
@@
We note that these metrics strongly depend on the progenitor’s extinction $A_{V}$ and background density. The latter is characterized by the number of background stars $N_{\rm bg}$ (accounting for the CMD selection in §II.2) within the $10^{\circ}$ search radius. In the upper row of Fig. 1, we show the completeness and purity as functions of these metrics. The values are directly taken from the with-extinction case of C25 (see §3.7 therein), excluding extremely high extinction cases following §II.2. As expected, both completeness and purity decrease with $A_{V}$ and $N_{\rm bg}$. They become less than 20% at $A_{V}>0.6$ or $N_{\rm bg}>6\times 10^{6}$. These values approximately correspond to the low-Galactic latitude region $|b|\lesssim 20^{\circ}$, where the high dust reddening and background contamination significantly affect the detection quality. On the other hand, the method achieves acceptable completeness and purity for the low-$A_{V}$ and low-$N_{\rm bg}$ GCs. For mock GCs with $A_{V}<0.6$ and $N_{\rm bg}<6\times 10^{6}$, the median completeness and purity are 50% and 59%, respectively. These values can even grow to $60\%-80\%$ near the Galactic poles.
@@ III.2 New streams
In Table 1, we list the key properties of the 34 streams in our high-quality sample. We provide information about the low-quality sample in Appendix C, Table 2. We also plot the high-quality sample in the great circle frame $\phi_{1}$$\phi_{2}$ in Fig. 2. The $\phi_{1}$$\phi_{2}$ frame is defined such that the progenitor GC is at $(0,0)$, and the stream is elongated along $\phi_{1}$. We achieve this by rotating the sky coordinates to minimize the standard deviation of $\phi_{2}$ for simulated tracer particles. We rank the streams in Fig. 2 by the length of simulated streams, which is characterized by the 90th percentile of angular separation, $r_{90}$. Since we only detect streams within the $10^{\circ}$ search radius around the GC, there are 17 streams with $r_{90}$ exceeding this radius. In §III.3, we perform a followup detection for these streams by extending the search radius to $20^{\circ}$.
@@
In Fig. 3, we show the CMD of detected stream stars using Gaia photometry. Most stars are gathered around or below the main sequence turn off. Since StarStream takes into account the color and magnitude uncertainties that typically grow with the $G$-magnitude, the spread around isochrone also becomes larger near the fainter end for detected stars. On the other hand, the spread of simulated streams is only due to the distance spread and is thus almost invariant of $G$-magnitude.
@@
Table 1: Summary of GC and stream properties for the high-quality sample. The Galactic longitude $l$, latitude $b$, heliocentric distance $d_{\odot}$, cluster mass $M_{\rm GC}$, and 3D half-mass radius $r_{\rm h}$ are taken from the M. Hilker et al. (2019) catalog. The extinction $A_{V}$ is either from D. Massari et al., W. E. Harris (1996) catalog, or D. J. Schlegel et al. (1998) map, see §II.2. $N_{\rm bg}$ is the number of CMD-selected Gaia DR3 stars within the $10^{\circ}$ search radius. The number of detections $N_{\rm detect}$, tidal frequency $\Omega_{\rm tid}$, and cluster mass loss rate $|\dot{M}|$ are calculated in this work. A plain ASCII version is available at https://github.com/ybillchen/StarStream_DR.
@@
III.3 Extended detection for long streams
@@
In Fig. 4, we show the two streams, NGC 5272 and 1851, with more than 10 new detections outside $10^{\circ}$. Note that the number of background stars approximately triples in the annulus, significantly reducing the signal-to-noise ratio in this region. Therefore, we should not expect the completeness and purity to stay unchanged in the extended area.
@@
III.4 Comparison with previous detections
@@
In addition to automated stream searches, there has also been detections of streams or tidal features around individual clusters, including five GCs by A. Sollima et al. (2019), four GCs by N. Shipp et al. (2018), NGC 288 (A. E. Piatti, 2018; S. Kaderali et al., 2019; C. J. Grillmair, 2025), NGC 1261 and 1904 (P. Awad et al., 2025), NGC 4590 (C. G. Palau & J. Miralda-Escudé, 2021), NGC 5024 and NGC 5053 (Z. Yuan et al., 2020), NGC 5272 (Y. Yang et al., 2023), NGC 5466 (J. Jensen et al., 2021), NGC 5904 (C. J. Grillmair, 2019), NGC 6341 (G. F. Thomas et al., 2020), and NGC 7089 (C. J. Grillmair, 2022). We do not present a detailed, stream-by-stream comparison here, as most of these detections are already in the R. Ibata et al. (2024) catalog. Moreover, many of these studies do not release data products that are directly comparable with our data, and they typically emphasize purity over completeness through strict selection criteria. However, we still find that the stream morphologies and kinematics reported in these works are consistent with our detections within the $10^{\circ}$ search radius.
Nevertheless, some individually detected streams are not in R. Ibata et al. (2024). For example, P. Awad et al. (2025) reported streams around NGC 1261 and NGC 1904 using the Southern Stellar Stream Spectroscopic Survey (S5; T. S. Li et al., 2019), and NGC 1904 is not in R. Ibata et al. (2024). By examining stream stars outside the tidal radius, we find that 53% (NGC 1261) and 27% (NGC 1904) of their reported stream stars are recovered in our catalog. Given that the completeness of our method is $\sim$60% for NGC 1261 and $\sim$30% for NGC 1904 (as indicated from Fig. 1), we can infer a purity $\sim$90% for their detections. On the other hand, our completeness is likely more than three times higher, as we detect an order of magnitude more stream stars around these clusters. This level of agreement is encouraging, especially given that we do not incorporate radial velocities or metallicity information from spectroscopic surveys.
Recently, P. B. Kuzma et al. (2025) identified potential tidal structures within $5^{\circ}$ around 22 MW GCs using the Pristine-Gaia synthetic catalog from Pristine Data Release 1 (E. Starkenburg et al., 2017; N. F. Martin et al., 2024). 12 of them overlap our high-quality sample: NGC 362, 1261, 1851, 1904, 5272, 6205, 6341, 6934, 6981, 7078, 7089, 7099. It is worth noting that many of their tidal structures do not show stream-like feature. This agrees with our finding that although many GC streams are “irregular”, the tidal streams around GCs are more common than was previously detected.
IV Mass loss rate of globular clusters
IV.1 Calculation of mass loss rate
@@
Y. Chen et al. (2025b) approximates $f_{\rm sel}$ as a Gaussian tube around the stream track when analyzing the R. Ibata et al. (2024) stream catalog. Here, however, we can make further simplification considering the quantified detection ratio $f_{\rm detect}$ between the expectation of the number of detections $N_{\rm detect}$ by StarStream and the true number $N_{\rm true}$ (see C25, ). In Appendix B, we prove that simply replacing the integral in Eq. (1) by the total number of tracer particles $N_{\rm tracer}$ times $f_{\rm detect}$ yields an unbiased estimate of the mass loss rate:
@@
where $f_{\rm detect}\equiv N_{\rm detect}/N_{\rm true}$ varies with individual streams. Using the tests by C25, we find that $f_{\rm detect}$ is always centered at 0.9 for the high-quality sample, with a log-normal scatter that depends on the background density $N_{\rm bg}$. In Appendix B, we approximate the standard deviation of the log-normal scatter $\sigma_{\log f}$ as a linear function of $\log_{10}N_{\rm bg}$: $\sigma_{\log f}$ increases from 0.15 dex to 0.5 dex when $\log_{10}N_{\rm bg}$ increases from 5.5 to 7, and is fixed to $\sigma_{\log f}=0.15$ dex below $\log_{10}N_{\rm bg}<5.5$ (see Fig. 6). Since the purity and completeness of the low-quality sample are both low with large variation, the scatter can be more than 1 dex. Therefore, we exclude the low-quality sample for the calculation of the mass loss.
Note that $M_{\rm sel}$ and $N_{\rm tracer}$ only account for tracer particles within the $10^{\circ}$ search radius. The detection ratio is not calibrated outside $10^{\circ}$ and is likely much lower. We thus exclude extended stream segments outside this radius (see §III.3) for the subsequent calculation.
@@
Since we focus on the $10^{\circ}$ cone around each GC, the heliocentric distance of each stream member $d_{\odot,i}$ is very close to the distance to the center GC, $d_{\odot,\rm GC}$. Using the simulated stream, we verify that the standard deviation of distances is typically $0.025$ dex (at most $0.08$ dex). Therefore, the distance spread contributes negligibly to the total uncertainty since the intrinsic scatter due to the detection method itself is already >0.15>0.15 dex (see Appendix B). It is thus reasonable to adopt $d_{\odot,i}\approx d_{\odot,\rm GC}$. Correspondingly, we can define $m_{{\rm min},i}=m_{\rm min}(d_{\odot,\rm GC})\equiv m_{\rm min,GC}$ and $w_{i}=w(m_{\rm min,GC})\equiv w_{\rm GC}$. Eq. (3) is thus simplified to
@@
We also note that several GCs in our catalog are associated with the Sagittarius (Sgr) stream (M. Bellazzini et al., 2020; E. Vasiliev et al., 2021), including NGC 6715 (the nucleus), Arp 2, Ter 7, Ter 8, Whiting 1, and Pal 12. Among these systems, Whiting 1 and Pal 12 are in our high-quality sample analyzed here. Our selection may include stars from the Sgr stream, whose spatial, proper motion, color, and magnitude distributions are similar to those of these GCs. As a result, the inferred number of stream stars may be overestimated, and we therefore interpret our derived $\dot{M}$ as upper limits for these clusters.
IV.2 Mass loss rate vs. other cluster properties
@@
Figure 5 shows the measured mass loss rate for the high-quality sample as functions of GC’s mass $M_{\rm GC}$, effective tidal frequency $\Omega_{\rm tid}$, and 3D half-mass radius $r_{\rm h}$. The values of $M_{\rm GC}$ and $r_{\rm h}$ are directly taken from the M. Hilker et al. (2019) catalog, while $\Omega_{\rm tid}$ is approximated by $\sqrt{2}$ times the orbital frequency, which is characterized by the time between two adjacent apocentric passages. This definition is consistent with M. Gieles & O. Y. Gnedin (2023), who used the singular isothermal sphere profile and adopted F. Renaud et al. (2011)’s definition based on eigenvalues of the tidal tensor. The orbital frequency is computed via integrating the GC’s orbit in our Galactic potential model. These values are listed in Table 1.
@@
$\begin{array}[]{rl}|\dot{M}_{\rm ref}|&=6.8^{+1.6}_{-1.3}\,{\rm M}_{\odot}\,{\rm Myr^{-1}}\\ \sigma_{\rm int}&=0.27\pm 0.06\ {\rm dex}\\ a&=0.09\pm 0.18\\ c&=-0.20\pm 0.39\end{array}$ (6)
@@
where the uncertainties are obtained from 1000 bootstrap resampling. We find that the $\dot{M}$$M_{\rm GC}$ slope $a$ is consistent with zero within the standard deviation, and the value lies between the no black holes (BHs) model ($a=1/3$) and BHs models (a=1/3a=-1/3, assuming the initial mass $M_{\rm i}=2\times 10^{5}\,{\rm M}_{\odot}$) of M. Gieles & O. Y. Gnedin (2023). Our $a$ is lower than that in Y. Chen et al. (2025b, $a=0.66\pm 0.37$) by $1.6\sigma$. The $\dot{M}$$r_{\rm h}$ slope is slightly below but still consistent with zero. It is also consistent with Y. Chen et al. (2025b, $c=0.12\pm 0.72$). We also observe larger intrinsic scatter compared to Y. Chen et al. (2025b). This is likely because our measurement includes the uncertainty of the detection method itself, whereas Y. Chen et al. (2025b) assumed zero uncertainty for the detection method. Although trends can differ, the amplitude of $\dot{M}$ in this work is consistent with M. Gieles & O. Y. Gnedin (2023) and Y. Chen et al. (2025b) within the intrinsic scatter across the full ranges of $M_{\rm GC}$, $\Omega_{\rm tid}$, and $r_{\rm h}$.
@@
In particular, we measured $|\dot{M}|\approx 10\ {\rm M_{\odot}\,Myr^{-1}}$ for Pal 5. This is higher than that in Y. Chen et al. (2025b, $|\dot{M}|\approx 3\ {\rm M_{\odot}\,Myr^{-1}}$) partially because of the 70% increase of $N_{\rm detect}$ (as mentioned in §III.4). However, this increase is insufficient to explain the 230% increase of $\dot{M}$. Since Y. Chen et al. (2025b) measured the average mass loss rate over the past $\sim 6$ Gyr for Pal 5 while this work only measures the past 1 Gyr, the different values likely suggest an accelerating mass loss rate for Pal 5. This scenario is consistent with the prediction by M. Gieles et al. (2021), who showed that Pal 5’s potential high BH abundance ($f_{\rm BH}\approx 20\%$) could accelerate the mass loss rate from the initial $5-10\ {\rm M_{\odot}\,Myr^{-1}}$ to $10-20\ {\rm M_{\odot}\,Myr^{-1}}$ near the end of its lifetime.
@@
We also notice several low-mass but high-$\dot{M}$ GCs similar to Pal 5, including NGC 7492, Pal 12, and Whiting 1. These GCs have $M_{\rm GC}\lesssim 2\times 10^{4}\,{\rm M}_{\odot}$ but half-mass radii above average, $r_{\rm h}\gtrsim 10$ pc, making them much “fluffier” than average GCs. Their high mass loss rates place them in closer agreement with M. Gieles & O. Y. Gnedin (2023)’s BHs models (a=1/3a=-1/3), while they are seemingly outlier of other models. M. Gieles et al. (2021) suggested that these GCs are also BH-rich and are reaching complete tidal dissolution. They also predicted higher-than-average mass loss rates for these GCs, $|\dot{M}|\approx 10\ {\rm M_{\odot}\,Myr^{-1}}$. Our measurements agree with their prediction, providing observational support for the BH-rich scenario.
V Summary and discussion